OpenStudy (anonymous):

9. Which value for x makes the sentence true? 3x - 3 = 15

OpenStudy (anonymous):

That's just another way of saying solve for x.

OpenStudy (anonymous):

I know but plz answer very important

OpenStudy (anonymous):

Let's see 3x-3 =15 add three to both sides 3x+3 -3=15+3 Simplify 3x=18 x=?

OpenStudy (anonymous):

Answer choice is: x=6 x=3 x=21 x=18

OpenStudy (anonymous):

Never mind anymore :'( it was x=21

OpenStudy (ghostgate):

Its easy. First of all: 3x - 3 = 15 right? Add three to both sides: 3x - 3 + 3 = 15 + 3 To get: 3x = 18 Divide: x = 18 / 3 Which equals: x = 6 Hope that helps!

OpenStudy (anonymous):

Tux ghost

OpenStudy (anonymous):

21 was not the answer 6 is !

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Thx every1

OpenStudy (ghostgate):

No problem.